OpenStudy (anonymous):

How do I find the axis of symmetry in the below functions? 3(x + 4)^2 + 1 and 2x^2 −16x + 15 Would the axis of symmetry in the first one be 4?

OpenStudy (anonymous):

or -4?

OpenStudy (amoodarya):

y=a(x-b)^2+c x=b is axis of symmetry

OpenStudy (anonymous):

How would I find the axis of symmetry in 2x^2 −16x + 15 though?

OpenStudy (anonymous):

-b/2a?

OpenStudy (amoodarya):

ok turn it to standard form

OpenStudy (anonymous):

It is in standard form.

OpenStudy (anonymous):

Would it be 16 / 4? Would the axis of symmetry in 2x^2 −16x + 15 be 4?

OpenStudy (anonymous):

And the axis of symmetry in 3(x + 4)^2 + 1 is -4 correct?

OpenStudy (amoodarya):

\[2x^2 −16x + 15 =2(x^2-8x) +15\] \[2(x^2 −8x) + 15=2((x-4)^2 -4^2 )+15\]

OpenStudy (amoodarya):

3(x + 4)^2 + 1 is -4 yes it is correct

OpenStudy (anonymous):

And the one in standard form is 4. Correct?

OpenStudy (cybershadow):

in the standard form it wud be x= -b/2a

OpenStudy (anonymous):

I pointed that out, but he put it in the y = form anyways xD

OpenStudy (cybershadow):

-(-16)/2x2 = 16/4 = 4 YES correct

OpenStudy (anonymous):

Yay, thanks!

OpenStudy (cybershadow):

u r on a roll tongiht haha well done!

OpenStudy (amoodarya):

yes for second x=4 is ok

OpenStudy (amoodarya):

"CyberShadow' I don't know that she knew derivation

OpenStudy (cybershadow):

haha no worries, I dont Care abt medalss xD lol

OpenStudy (cybershadow):

i got ur point Sir @amoodarya nice explanation :)